OpenStudy (anonymous):

C6H12H6 --> 2C2H5OH + CO2 Find the mass of C2H5OH from 750g of C6H12O6

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@e.mccormick @e.cociuba

OpenStudy (abb0t):

So, you have your starting product: \(\sf \color{red}{750~g}\) of glucose. Now, first, make sure that your equation is balanced. Always start with the given information first. So, you take the \(\sf \color{red}{750~g}\) of Glucose \(\times\) molar mass of \(\sf \color{purple}{C_6 H_{12} O_6}\) this should cancel out units to get moles! Remember that molar mass has units of \(\sf \frac{grams}{mol}\) so, you're dividing!

OpenStudy (abb0t):

Once you have moles, note that you have a 1:2 ratio. Meaning, to get from moles of glucose to moles of \(\sf \color{red}{C_2H_5OH}\) you must multiply it by 2! So, \(\sf mol~of~glucose \times\frac{2~mol~C_2H_5OH}{1~mol~C_6H_{12}O_6}\) THEN you can finally find the mass by multiplying again by the molar mass of \(\sf C_2H_5OH \)

OpenStudy (abb0t):

The units should cancel out.